OpenStudy (anonymous):

What is the degree of the polynomial 2st^4 + s^2t^2 - 9s^5t + 21 ? A) 2 B) 4 C) 6 D) 21

OpenStudy (anonymous):

What is the degree of the polynomial t^2 + 12t + 36? A) 1 B) 2 C) 3 D) 36

OpenStudy (anonymous):

here you have two variables so you have to consider the sum of their powers and see which one is the highest

OpenStudy (anonymous):

I don't know how to do these two... At all.

OpenStudy (anonymous):

you have two variables s and t

OpenStudy (anonymous):

in each term check their respective powers and add them

OpenStudy (anonymous):

whichever sum is the highest, that is the degree...

OpenStudy (anonymous):

So, what answer have you come up with???

OpenStudy (anonymous):

@TuringTest Thanks mate ☺

OpenStudy (anonymous):

I haven't... I'm still confused. I don't know how to add it all.

OpenStudy (anonymous):

2st^4 - power of s is 1 and of t is 4 so total 1+4 = 5 s^2t^2 - power of s is 2 and of t is 2 so total 2+2 = 4 9s^5t - power of s is 5 and of t is 1 so total 5+1 = 6 21 - power is zero so now what is your answer???

OpenStudy (anonymous):

um... 6?

OpenStudy (anonymous):

correct......☺

OpenStudy (anonymous):

Yay:) So what about #2?

OpenStudy (anonymous):

In second one there is only one variable t so can't you see what is its highest power/index???

OpenStudy (anonymous):

so 2? B

OpenStudy (anonymous):

Correct...☺

OpenStudy (anonymous):

yay:D thank you so much !